OpenStudy (anthonyn2121):

Help please! Use the Mean Value Theorem to 1) Find the point c on (a,b) at which the tangent line of the following function is parallel to the secant line 2) Write an equation for the tangent line f(x)=-x^2+6x [1,4]

myininaya (myininaya):

so you know you need this: \[f'(c)=\frac{f(b)-f(a)}{b-a}\]

myininaya (myininaya):

f'(c) tells you to find the derivative of f and then evaluate f' at c -- b tells you to plug in 4 a tells you to plug in 1

OpenStudy (anthonyn2121):

Okay I got f'(x) -2x+6 and [f(4)-f(1)]/3 = 1. And x=7/2

myininaya (myininaya):

so -2x+6=1 -2x=1-6 right

myininaya (myininaya):

1-6 isn't -7 but -5 right?

myininaya (myininaya):

i think you know that -2x=-5 so x=5/2 not 7/2

myininaya (myininaya):

so you are asked to find the tangent line to the point (5/2,f(5/2)) to the curve f=-x^2+6x

myininaya (myininaya):

so you do this: y-f(5/2)=f'(5/2)*(x-5/2) you need to find f(5/2) and also f'(5/2)

OpenStudy (anthonyn2121):

I got f'(5/2)=1 and f(5/2)=35/4. So my final answer is y-(35/4)=x-(5/2)

myininaya (myininaya):

looks great

OpenStudy (anthonyn2121):

Thank you so much

myininaya (myininaya):

np you did all the work